OpenStudy (anonymous):

x – 2y < 18 4x + y > –5 Explain, in complete sentences, how you would use the graphing method to solve the following system of inequalities and what the final solution will look like on the coordinate plane.

OpenStudy (anonymous):

Just use Geogebra and describe it... o.O. and just use the slope and count rise over run and stuff o.o. substitute in what y equals and......color in the part that's lower than the part that isn't on 18 in the 1st graph or -5 U_U

OpenStudy (cwrw238):

first draw the graphs of the 2 lines x - 2y = 18 and 4x + y = -5 plot 3 points for each line

OpenStudy (anonymous):

how would i find the points

OpenStudy (cwrw238):

just make them up eg let x = 0 then for x - 2y = 18, 0 - 2y = 18 , y = 18/-2 = -9 so plot the point (0,-9) set up a table for each line

OpenStudy (anonymous):

thankyou!

OpenStudy (cwrw238):



OpenStudy (cwrw238):

it will look something like this